# The differences of baryonic and dark matter scaling relations in galaxy clusters between IllustrisTNG simulations and observations

### v1

@@ 5 Discussions
5.1 Mocks and the fraction of baryons
In an observational context, the measurement of masses is limited by the two-dimensional (2D) projection of the sky, which introduces biases in the estimates. Among the methods used to estimate mass, the Sunyaev-Zel’dovich (SZ) effect is useful for calculating $M_{\rm 500c}$, while X-ray $\beta$-model fitting is useful for calculating $M_{\rm hot-gas,500c}$. In contrast, simulations tend to overestimate observable values when using three-dimensional (3D) projections instead of 2D projections. Nevertheless, simulations can generate mocks that provide an estimate of the mean bias, $b$, between the mass computed in simulations, $M_{\rm 500c}^{\rm SIM}$, and the observable mass in mocks, $M_{\rm 500c}^{\rm EST}$ (Pop et al., 2022b).
The relationship between these quantities is defined as,
@@ 5.1 Mocks and the fraction of baryons
If we consider both values for the estimation of the total mass $M_{\rm 500c}$, defined as $M_{\rm 500c}^{\rm EST}=(1-b_{\rm i})M_{\rm 500c}^{\rm SIM}$, with $b_{\rm i}=\{b_{\rm X},b_{\rm SZ}\}$, we define $f_{\rm bar,500c}^{\rm EST}\equiv M_{\rm bar,500c}/M_{\rm 500c}^{\rm EST}$ as the synthetic baryon fraction, and $f_{\rm bar,500c}^{\rm SIM}$ as the simulated value. The relation between both baryon fractions can be written as $f_{\rm bar,500c}^{\rm EST}\equiv(1-b_{f})f_{\rm bar,500c}^{\rm SIM}$, where $b_{f}$ is the baryon fraction bias. The parameter $b_{f}$ can be related to $b_{\rm i}$, and its associated uncertainty $\sigma_{b_{f}}$, through the following expressions,
$b_{f}\equiv 1-\frac{1}{1-b_{\rm i}}$ (11)
$\sigma_{b_{f}}=\frac{\sigma_{b_{\rm i}}}{\left(1-b_{\rm i}\right)^{2}}$ (12)
Using the values from Pop et al. (2022b), we find $b_{f}\approx-0.1429\pm 0.0039$ (for $b_{\rm SZ}$) and $b_{f}\approx-0.2048\pm 0.0058$ (for $b_{\rm X}$). Under these assumptions, we expect an overestimation in the baryon fraction from 14.29 to 20.48 per cent, which represents a significant departure from the results of C18. However, this overestimation could be reduced if the baryonic mass bias $b_{\rm bar}$ could be estimated. A condition of $b_{\rm bar}>b_{\rm i}$ would be required to produce a decrease in $f_{\rm bar,500c}^{\rm EST}$ relative to $f_{\rm bar,500c}^{\rm SIM}$. Moreover, $b_{\rm bar}$ could account for individual biases associated with the stellar mass and the various gas phases.
@@ 5.2 Mass accretion in FoF masses
It is worth noting that the FoF clustering algorithm assigns all particles gravitationally bound to the same halo. Because of this, this approach does not recover the universal baryon fraction, whereas methods using large concentric radii can achieve this at $1.5-2.5R_{\rm 200c}$ (Ayromlou et al., 2023). The concentric radius method, however, introduces its own challenges. Specifically, it may include elements not gravitationally bound to the cluster and does not account for the cluster’s shape, as it only considers elements within a sphere. It would be particularly interesting to investigate the relationship between the shape of galaxy clusters and their baryon fraction using the concentric radius method.
Returning to the FoF algorithm, the differences between the baryon fraction reported by Planck Collaboration et al. (2016a), and that of IllustrisTNG haloes could originate from two main explanations:
  • The galaxy clusters are accreting dark matter efficiently.
  • The galaxy clusters are accreting gas inefficiently.
@@
Both hypotheses are related to the mass accretion rate (MAR) of dark matter ($\rm MAR_{dm}$) being higher than that of gas ($\rm MAR_{gas}$). Pizzardo et al. (2023b) developed a caustic technique to measure the mass accretion rate of galaxy clusters using spherical shells. The radii of these spherical shells are determined by the radial velocity profile, where the radial velocity satisfies $v_{\rm rad}\leq A\times v_{\rm min}$, with $v_{\rm min}$ being the minimum radial velocity in the cluster. $A$ is an empirically determined adjustment parameter, which is chosen to accurately capture the infall region described causticly. Pizzardo et al. (2023b) selected $A=0.72$ to have a $r_{3c}={\rm MAR_{3D}/MAR_{c}}=1.00\pm 0.04$, where $\rm MAR_{3D}$ and $\rm MAR_{C}$ are mass accretion rates obtained by 3D projections and by the caustic description respectively. Using TNG300, Pizzardo et al. (2023b) analyzed 1697 FoF haloes with $M_{\rm 200c}\geq 10^{14}\,\rm M_{\odot}$, covering a redshift range of $0.01\leq z\leq 1.04$, ultimately working with 1318 haloes. The remaining $\sim$22 per cent haloes were excluded due to limitations in the robust application of the caustic technique (Pizzardo et al., 2023a). An interesting result pertains to the mass accretion rates of galaxies and dark matter within these haloes. The study found that the $\rm MAR_{dms}$ are approximately 6.5 per cent lower than the galaxy mass accretion rates ($\rm MAR_{galaxies}$) (Pizzardo et al., 2023b). This discrepancy arises because the clustering amplitude of galaxies relative to dark matter is larger on smaller scales and at higher redshifts (Davis & Peebles, 1983; Davis et al., 1985; Jenkins et al., 1998).
@@
However, a challenge arises when attempting to apply this model to gas. This difficulty stems from the fact that IllustrisTNG employs a cell-based method to track the evolution of gas (Pillepich et al., 2018). With this approach, tracing the gas becomes challenging as its vectorial information is not retained. Consequently, a gas particle simulation suite is required for more accurate tracking. For instance, the FLAMINGO simulations (Schaye et al., 2023; Kugel et al., 2023) utilize a gas particle method in simulation boxes of 1.0 and 2.8 Gpc. The comparison of these results with FLAMINGO simulations, as well as the tracking of gas and its mass accretion rate, will be addressed in a future study.
@@ 6 Summary and conclusions
In this study, we examined the scaling relation differences between the state-of-the-art magnetohydrodynamic and cosmological simulations of IllustrisTNG and the 91 galaxy clusters identified via the SZ effect in the SPT-SZ survey (Bleem et al., 2015), as reported by C18. Our analysis focused on a simulation sample consisting of 218 haloes from TNG100 and 1604 haloes from TNG300, each with $M_{\rm 200c}\geq 7\times 10^{13}\,\rm M_{\odot}$. The baryonic mass fraction was defined as a function of halo mass within a concentric sphere of radius $R_{500}$, using the subfind algorithm to select galaxy components.
@@
We defined two observational comparison cases: (1) incorporating only the hot gas in the halo and galaxy stars within a spherical aperture of $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$, and (2) including all baryonic components (gas and stars) in the halo, both analyzed at $R_{\rm 500c}$. To test the self-similarity of the haloes, we applied a SPL model and found that the slopes in both cases and simulations converge towards $\alpha_{\rm SPL}\to 0$, indicating the elimination of redshift dependence in the scaling relations.
Our main findings can be summarized as follows:
@@
  • The slopes of the TNG100 and TNG300 haloes in the SPL scaling relations ($\alpha_{\rm SPL}$) align with the scaling relation slope values reported by C18 within error of $1-2\sigma$ for the ”hot-gas + $r_{\rm op}$ stars” case. This suggests a potential solution to the missing baryons problem identified in C18 and the tensions within the hierarchical formation scenario.
@@
  • When including all baryonic components (hot-gas, warm-gas, cold-gas, ICL, galaxy stars), the SPL scaling relations exhibit a null slope. Low-mass haloes ($\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\lesssim 14$) show significantly elevated baryonic fractions but do not reach the universal baryon fraction, whereas high-mass haloes ($\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\gtrsim 14$) achieve this value. This mass threshold correlates with the loss of self-similarity in the haloes and suggests that non-gravitational processes play a critical role in baryon redistribution.
@@
  • The mean percentage of missing baryons is notably higher in low-mass haloes ($\sim$17.67 per cent in TNG100, and $\sim$17.80 per cent in TNG300) compared to high-mass haloes ($\sim$6.95 per cent in TNG100, and $\sim$5.90 per cent in TNG300). This disparity accounts for the observed behavioral differences in scaling relations across different baryonic component configurations. Furthermore, the missing baryons in the IllustrisTNG haloes are primarily due to the exclusion of ICL and warm gas, particularly in low-mass haloes, while cold gas does not significantly contribute to the total missing components. This observation strongly supports the warm-hot intergalactic medium (WHIM) scenario.
@@
  • Our analysis indicates that the discrepancies in galaxy apertures, particularly in the characterization of ICL, can result in a significant variation of approximately 5 per cent in the percentage of missing baryons, depending on the method used to measure stellar mass. This variability must be considered in observational studies.
@@
We discussed the accretion of baryonic components within FoF haloes, highlighting that approximately $\sim$21 per cent of baryons are missing compared to the expected universal fraction (Planck Collaboration et al., 2016a). While the redistribution of baryons to the outskirts of low-mass haloes (Gouin et al., 2022; Ayromlou et al., 2023) may account for this deficit, it does not explain the missing baryons in high-mass haloes. We hypothesize that FoF haloes may either be efficiently accreting dark matter or inefficiently accreting gas. Although the dark matter mass accretion rate ($\rm MAR_{dm}$) can be traced in IllustrisTNG using a caustic technique (Pizzardo et al., 2023b, a), the gas mass accretion rate ($\rm MAR_{gas}$) cannot be similarly traced due to the absence of positional and vectorial information for gas cells in IllustrisTNG. Future work will focus on following the gas particle of SPH simulation suites, such as FLAMINGO, to track gas accretion more effectively and calculate $\rm MAR_{gas}$.
@@
Finally, we emphasized the importance of incorporating non-gravitational processes into scaling relation analyses, particularly through the BPL and GDPL models. These models allowed us to identify critical mass break points where self-similarity is either lost or preserved. Future observational surveys could validate and refine these models, especially when compared with current and upcoming simulations of large haloes. Such efforts may expand this framework by integrating additional data, further improving our understanding of baryonic scaling relations and their implications for cosmology.
@@ Appendix A Calculating baryonic component percentages
Figure 6: Percentage distributions of cold-gas ($T<10^{5}$ K in upper row panels), warm-gas ($10^{5}\leq T<10^{7}$ K in upper-center row panels), hot-gas ($T\geq 10^{7}$ K in bottom-center row panels) and all-gas (in bottom row panels) masses considering the baryonic masses. The histograms are divided in the haloes with $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}<14$ (left colum panels), $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}\geq 14$ (center colum panels) and all halo (right colum panels) masses. (Black histograms) TNG100 haloes. (Red histograms) TNG300 haloes. The dashed lines represent the normal distribution fitted in their respective histogram (with their respective color).
@@
Figure 7: Porcentual distributions stellar masses considering the baryonic masses. Stellar galaxy mass in $r_{\rm op}$ aperture (upper row panels), ICL mass in $r_{\rm op}$ aperture (upper center row panels), stellar galaxy mass in subfind aperture (bottom center row panels) and ICL mass in subfind aperture (bottom row panels). The histograms are divided in the haloes with $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}<14$ (left colum panels), $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}\geq 14$ (center colum panels) and all halo (right colum panels) masses. (Black histograms) TNG100 haloes. (Red histograms) TNG300 haloes. The dashed lines represent the normal distribution fitted in their respective histogram (with their respective color).
@@
Fig. 6 shows the distributions of the gas component percentages (cold, warm, and hot), while Fig. 7 illustrates the distributions of the stellar component percentages (galaxies and ICL, using $r_{\rm op}$ and subfind apertures). In this context, comparing these percentages with the total baryonic mass of the clusters is crucial, as it provides a straightforward way to analyze the scatter in TNG100 and TNG300 haloes. For instance, low-mass haloes exhibit a high scatter in the hot gas percentages, which consequently results in a high scatter across all haloes for this component. A similar trend is observed for warm and cold gas components, where the standard deviation is comparable to the mean values. This occurs due to the significant number of clusters with low percentages for these components, which affects the shape of the distribution. Nevertheless, there are numerous clusters with high warm and cold gas percentages, supporting the inclusion of these components as ”missing baryons” within the simulation environment. Additionally, the standard error, calculated as $\sigma/\sqrt{N}$ (where $N$ is the sample size), is on the order of tenths (see Table 3), making these percentages comparable to each other. Another baryonic component identified as missing is the ICL, which is derived from the difference between the stellar halo and the galaxies within the halo. The differences between $r_{\rm op}$ and subfind apertures are evident, with the latter capturing most of the stellar components of the haloes (see Section 4.1 and Fig. 1) for details.

### v2

@@ 5 Discussions
5.1 Mock observations and the fraction of baryons
In observational studies, the measurement of masses is limited by the 2D projection of the sky, which introduces biases into the estimates. Among the methods used to estimate mass, the SZ effect is useful to calculate $M_{\rm 500c}$, while the fitting of the X-ray model $\beta$ is useful for calculating $M_{\rm hot-gas,500c}$. In contrast, simulations often overestimate observable values when using 3D projections instead of 2D projections. However, simulations can generate mock observations that provide an estimate of the mean bias, $b$, between the mass computed in simulations, $M_{\rm 500c}^{\rm SIM}$, and the observable mass in mocks, $M_{\rm 500c}^{\rm EST}$ (Pop et al., 2022b). The relationship between these quantities is defined as
@@ 5.1 Mock observations and the fraction of baryons
If we consider both values for the estimation of the total mass $M_{\rm 500c}$, defined as $M_{\rm 500c}^{\rm EST}=(1-b_{\rm i})M_{\rm 500c}^{\rm SIM}$, with $b_{\rm i}=\{b_{\rm X},b_{\rm SZ}\}$, we define $f_{\rm bar,500c}^{\rm EST}\equiv M_{\rm bar,500c}/M_{\rm 500c}^{\rm EST}$ as the synthetic baryon fraction and $f_{\rm bar,500c}^{\rm SIM}$ as the simulated value. The relation between both baryon fractions can be written as $f_{\rm bar,500c}^{\rm EST}\equiv(1-b_{f})f_{\rm bar,500c}^{\rm SIM}$, where $b_{f}$ is the baryon fraction bias. The parameter $b_{f}$ can be related to $b_{\rm i}$, and its associated uncertainty $\sigma_{b_{f}}$, through the following expressions:
$b_{f}\equiv 1-\frac{1}{1-b_{\rm i}},$ (11)
$\sigma_{b_{f}}=\frac{\sigma_{b_{\rm i}}}{\left(1-b_{\rm i}\right)^{2}}.$ (12)
Using the values from Pop et al. (2022b), we find $b_{f}\approx-0.1429\pm 0.0039$ (for $b_{\rm SZ}$) and $b_{f}\approx-0.2048\pm 0.0058$ (for $b_{\rm X}$). Under these assumptions, we expect an overestimation in the baryon fraction from 14.29 to 20.48 per cent, which represents a significant departure from the results of C18. However, this overestimation could be reduced if the baryonic mass bias, $b_{\rm bar}$, could be estimated. A condition of $b_{\rm bar}>b_{\rm i}$ would be required to produce a decrease in $f_{\rm bar,500c}^{\rm EST}$ relative to $f_{\rm bar,500c}^{\rm SIM}$. Moreover, $b_{\rm bar}$ could account for individual biases associated with the stellar mass and the various gas phases.
@@ 5.2 Mass accretion in FoF masses
The FoF clustering algorithm assigns all gravitationally bound particles to the same halo. Consequently, this approach does not recover the universal baryon fraction,. In contrast, methods that use large concentric radii can achieve this at $1.5-2.5R_{\rm 200c}$ (Ayromlou et al., 2023). However, this method also presents challenges: it may include elements not gravitationally bound to the cluster and does not account for the cluster’s shape, as it only considers elements within a sphere. It would be particularly interesting to investigate the relationship between the shape of galaxy clusters and their baryon fraction using the concentric radius method.
The differences between the baryon fraction reported by Planck Collaboration et al. (2016a), and that found in IllustrisTNG haloes when using the FoF algorithm may be attributed to two main factors. First, galaxy clusters may be accreting dark matter efficiently. Second, they may be accreting gas inefficiently.
@@
Both hypotheses are related to the mass accretion rate (MAR) of dark matter ($\rm MAR_{dm}$) being higher than that of gas ($\rm MAR_{gas}$). Pizzardo et al. (2023b) developed a caustic technique to measure the mass accretion rate of galaxy clusters using spherical shells. The radii of these spherical shells are determined by the radial velocity profile, where the radial velocity satisfies $v_{\rm rad}\leq A\times v_{\rm min}$, with $v_{\rm min}$ being the minimum radial velocity in the cluster. The parameter $A$ is an empirically determined adjustment parameter, chosen to accurately capture the infall region described causticly. Pizzardo et al. (2023b) selected $A=0.72$ to have a $r_{3c}={\rm MAR_{3D}/MAR_{c}}=1.00\pm 0.04$, where $\rm MAR_{3D}$ and $\rm MAR_{C}$ are mass accretion rates obtained by 3D projections and caustic description, respectively. Using TNG300, Pizzardo et al. (2023b) analysed 1697 FoF haloes with $M_{\rm 200c}\geq 10^{14}\,\rm M_{\odot}$, covering a redshift range of $0.01\leq z\leq 1.04$, ultimately working with 1318 haloes. The remaining $\sim$22 per cent of the haloes were excluded due to limitations in the reliable application of the caustic technique (Pizzardo et al., 2023a). A notable finding concerns the mass accretion rates of galaxies and dark matter within these haloes. The study found that the $\rm MAR_{dms}$ are approximately 6.5 per cent lower than the galaxy mass accretion rates ($\rm MAR_{galaxies}$) (Pizzardo et al., 2023b). This discrepancy arises because the clustering amplitude of galaxies relative to dark matter is larger on smaller scales and at higher redshifts (Davis & Peebles, 1983; Davis et al., 1985; Jenkins et al., 1998).
@@
However, applying this model to gas presents additional challenges. In IllustrisTNG, the evolution of gas is tracked using a cell-based method(Pillepich et al., 2018). With this approach, tracing the gas becomes challenging as its vectorial information is not retained. Consequently, a gas particle simulation suite is required for more accurate tracking. The FLAMINGO simulations (Schaye et al., 2023; Kugel et al., 2023), for example, use a gas particle method in simulation boxes of 1.0 and 2.8 Gpc. A comparison of these results with FLAMINGO simulations, including an analysis of gas dynamics and mass accretion rate will be addressed in a future study.
@@ 6 Summary and conclusions
In this study, we examined the scaling relation differences between the state-of-the-art magnetohydrodynamic and cosmological simulations from IllustrisTNG and the 91 galaxy clusters identified via the SZ effect in the SPT-SZ survey (Bleem et al., 2015), as reported by C18. Our simulation sample consisted of 218 haloes from TNG100 and 1604 haloes from TNG300, each with $M_{\rm 200c}\geq 7\times 10^{13}\,\rm M_{\odot}$. The baryonic mass fraction was defined as a function of halo mass within a concentric sphere of radius $R_{500}$, with galaxy components selected using the Subfind algorithm.
@@
We defined two observational comparison cases: (1) incorporating only the hot gas within the halo and galaxy stars within a spherical aperture of $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$; and (2) including all baryonic components (gas and stars) in the halo, both analysed at $R_{\rm 500c}$. To test the self-similarity of the haloes, we applied an SPL model and found that the slopes in both cases and simulations converge towards $\alpha_{\rm SPL}\to 0$, indicating the elimination of redshift dependence in the scaling relations.
Our main findings can be summarised as follows.
@@
  • The slopes of the SPL scaling relations ($\alpha_{\rm SPL}$) in TNG100 and TNG300 haloes align with those reported by C18 within an error of $1-2\sigma$ for the ”hot gas + $r_{\rm op}$ stars” case. This suggests a potential solution to the missing baryons problem identified in C18 and the tensions within the hierarchical formation scenario.
@@
  • When all baryonic components (hot gas, warm gas, cold gas, ICL, and galaxy stars) are included, the SPL scaling relations exhibit a null slope. Low-mass haloes ($\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\lesssim 14$) show significantly elevated baryonic fractions, but remain below the universal baryon fraction, whereas high-mass haloes ($\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\gtrsim 14$) achieve this value. This mass threshold correlates with the loss of self-similarity in the haloes and suggests that non-gravitational processes play a critical role in baryon redistribution.
@@
  • The mean percentage of missing baryons is higher in low-mass haloes ($\sim$17.67 per cent in TNG100, $\sim$17.80 per cent in TNG300) compared to high-mass haloes ($\sim$6.95 per cent in TNG100, $\sim$5.90 per cent in TNG300). This disparity accounts for the observed behavioural differences in scaling relations across different baryonic component configurations. Furthermore, the missing baryons in the IllustrisTNG haloes are primarily due to the exclusion of ICL and warm gas, particularly in low-mass haloes, whereas cold gas contributes negligibly. This finding strongly supports the warm-hot intergalactic medium (WHIM) scenario.
@@
  • We find that discrepancies in galaxy apertures, particularly in the characterisation of ICL, can lead to variation of approximately five per cent in the inferred missing baryon fraction, depending on the stellar mass measurement model. This variability must be considered in observational studies.
@@
We investigated the accretion of baryonic components within FoF haloes and found that approximately $\sim$21 per cent of baryons are missing compared to the expected universal fraction (Planck Collaboration et al., 2016a). While the redistribution of baryons to the outskirts of low-mass haloes (Gouin et al., 2022; Ayromlou et al., 2023) may account for this deficit, it does not explain the shortfall in high-mass haloes. We hypothesise that FoF haloes may either be efficiently accreting dark matter or inefficiently accreting gas. Although the dark matter mass accretion rate ($\rm MAR_{dm}$) can be traced in IllustrisTNG using a caustic technique (Pizzardo et al., 2023b, a), the gas mass accretion rate ($\rm MAR_{gas}$) cannot be similarly traced due to the absence of positional and vectorial information for gas cells. Future work will focus on following gas particles in SPH simulation suites, such as FLAMINGO, to track gas accretion more effectively and calculate $\rm MAR_{gas}$.
@@
Finally, we emphasised the importance of incorporating non-gravitational processes in scaling relation analyses, particularly through the BPL and GDPL models. These models enable the identification of critical mass breakpoints where self-similarity is either lost or preserved. Future observational surveys could validate and refine these models, particularly when compared with current and upcoming simulations of large haloes. Such efforts may expand this framework by integrating additional data, further improving our understanding of baryonic scaling relations and their implications for cosmology.
@@ Appendix A Calculating baryonic component percentages
Figure 6: Percentage distributions of cold-gas ($T<10^{5}$ K in upper row panels), warm-gas ($10^{5}\leq T<10^{7}$ K in upper-center row panels), hot-gas ($T\geq 10^{7}$ K in bottom-center row panels) and all-gas (in bottom row panels) masses considering the baryonic masses. The histograms are divided in the haloes with $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}<14$ (left colum panels), $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}\geq 14$ (center colum panels) and all halo (right colum panels) masses. (Black histograms) TNG100 haloes. (Red histograms) TNG300 haloes. The dashed lines represent the normal distribution fitted in their respective histogram (with their respective colour).
@@
Figure 7: Porcentual distributions stellar masses considering the baryonic masses. Stellar galaxy mass in $r_{\rm op}$ aperture (upper row panels), ICL mass in $r_{\rm op}$ aperture (upper centre row panels), stellar galaxy mass in Subfind aperture (bottom centre row panels) and ICL mass in Subfind aperture (bottom row panels). The histograms are divided in the haloes with $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}<14$ (left colum panels), $\log_{10}{\left(M_{\rm 500c}/{\rm M_{\odot}}\right)}\geq 14$ (centre colum panels) and all halo (right colum panels) masses. (Black histograms) TNG100 haloes. (Red histograms) TNG300 haloes. The dashed lines represent the normal distribution fitted in their respective histogram (with their respective colour).
@@
Fig. 6 shows the distributions of the gas component percentages (cold, warm, and hot), while Fig. 7 illustrates the distributions of the stellar component percentages (galaxies and ICL, using $r_{\rm op}$ and Subfind apertures). In this context, comparing these percentages with the total baryonic mass of the clusters is crucial, as it provides a straightforward way to analyse the scatter in TNG100 and TNG300 haloes. For instance, low-mass haloes exhibit a high scatter in the hot gas percentages, which consequently results in a high scatter across all haloes for this component. A similar trend is observed for warm and cold gas components, where the standard deviation is comparable to the mean values. This occurs due to the significant number of clusters with low percentages for these components, which affects the shape of the distribution. Nevertheless, there are numerous clusters with high warm and cold gas percentages, supporting the inclusion of these components as ”missing baryons” within the simulation environment. Additionally, the standard error, calculated as $\sigma/\sqrt{N}$ (where $N$ is the sample size), is on the order of tenths (see Table 3), making these percentages comparable to each other. Another baryonic component identified as missing is the ICL, which is derived from the difference between the stellar halo and the galaxies within the halo. The differences between $r_{\rm op}$ and Subfind apertures are evident, with the latter capturing most of the stellar components of the haloes (see Section 4.1 and Fig. 1) for details.